# Uncertainty Quantification on Graph Learning: A Survey

### v2

@@ V-B1 Conformal Prediction Overview
Inductive CP, or called split conformal prediction (SCP), is a widely applied version of CP. For a trained classifier ${f}$ and input $x$ with label space $\{1,2,...,K\}$, the estimated probability for each class is $\hat{y}=f(x)\in[0,1]^{K}$. Let $f(x)_{y}$ denote the true class from the final softmax layer. The calibration set $D_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is applied to quantify how uncertain the model $f$ would be. The score function $s(x,y)=1-f(x)_{y}$ is widely selected for classification tasks, and it measures how predictions conform to the true values at inputs. The conformal score for $(X_{i},Y_{i})\in D_{cal}$ is defined as $s_{i}=1-f(X_{i})_{Y_{i}}$. Let $S=\{s_{i}\}_{i=1}^{n}$, the empirical quantile of $S$ is given by
$Q_{1-\alpha}(S)={\lceil(1-\alpha)(n+1)\rceil}/{n}\text{-th quantile of }S.$ (28)
@@
If a test sample $(X_{n+1},Y_{n+1})$ drawn exchangeably with the elements of $D_{cal}$, we apply CP to construct a prediction set ${C}(X_{n+1})=\{k|1-f(X_{n+1})_{k}\leq Q_{1-\alpha}\}$, which includes all classes $k\in\{1,...,K\}$ whose conformal score is less or equal to $Q_{1-\alpha}$. Under the assumption of exchangeability, the coverage guarantee holds that
@@
Transductive CP, also referred to as full conformal prediction (FCP), is the first proposed method for conducting CP [111]. With training data $D_{train}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a test input $X_{n+1}$, the trained model $f^{k}$ is based on $D_{train}\cup\{(X_{n+1},k)\},\forall k\in\{1,2,...,K\}$, so multiple models need to be trained. For each $f^{k}$, the conformal scores are $s_{i}^{k}=1-f^{k}(X_{i})_{Y_{i}}$ and $s_{n+1}^{k}=1-f^{k}(X_{n+1})_{k}$. If $s_{n+1}^{k}$ is less than the $1-\alpha$ quantile of $\{s_{i}^{k}\}_{i=1}^{n}$, we put the corresponding $k$ to the prediction set of $X_{n+1}$.
@@
To make the size of $C(X_{n+1})$ more adaptive to $X_{n+1}$ under the exchangeability assumption, novel conformal score functions are proposed, such as  [112, 113, 114, 115]. However, the exchangeability assumption can be violated by distribution shift between calibration distribution $P_{XY}$ and test distribution $Q_{XY}$ such that $(X_{n+1},Y_{n+1})$ is non-exchangeable with the elements of $D_{cal}$. In the case of covariant shift ($P_{X}\neq Q_{X}$), [116] introduces an importance weighting technique based on likelihood ratio to maintain the coverage guarantee in Eq.(29). When concept shift ($P_{Y|X}\neq Q_{Y|X}$) occurs, the authors of [117, 118] focus on keeping $1-\alpha$ coverage on test data under dynamic shift with time, and the authors of [119, 120, 121] ensure the coverage guarantee on test data under stationary distribution shift (does not change with time).  [122] bounds the coverage gap by the total variation distance between calibration and test conformal scores under the non-exchangeability condition.
V-B2 Conformal Prediction for Graphs
For node classification tasks on graphs, the exchangeability may be violated for some specific settings [123]. First, in the transductive settings, the model has access to the entire graph structure and all the node features during training, calibration, and testing, but the labels of calibration and testing set are not available during training. In this case, the union of calibration and the unlabeled sample is exchangeable. However, in the inductive settings, the model only has access to the subgraph induced by the nodes in the training set $\mathcal{D}_{train}$ during training. The assumption of exchangeability does not hold in this case.
The authors in [124] adopt the jackknife method to estimate the uncertainty of a node $X_{i}$ inferred by a GNN. Specifically, uncertainty $\mathbb{U}(X_{i})=\mathbb{C}^{+}(X_{i})-\mathbb{C}^{-}(X_{i})$ is the difference between the lower and upper bound of the predictive confidence interval. Both bounds are defined by leave-one-out (LOO) loss, which are approximated by the influence function.
Another branch of research proposes new score functions for CP on graphs. The diffused adaptive prediction sets (DAPS) is introduced in [123]. DAPS exploits the graph structure based on neighborhood diffusion, and defines the diffused score as
$\hat{s}(X_{i},Y_{i})=(1-\lambda)s(X_{i},Y_{i})+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{X_{j}\in\mathcal{N}_{i}}s(X_{j},Y_{i}),$ (30)
where $\lambda$ signifies the diffusion parameter. Note that the diffused score function preserves exchangeability if the original scores are exchangeable. This diffusion process proves particularly advantageous in homophilous graphs, where connected nodes exhibit similar ground truth distributions. The rationale behind this is that for a node surrounded by mostly unperturbed neighbors, its probability can be better estimated using its neighbors’ information, as long as these vectors are sufficiently similar.
For the non-exchangeable conformal prediction cases, the procedure assumes a choice of deterministic fixed weights $w_{1},\dots,w_{n}\in[0,1]$ [122]. Then the prediction set concerning the weighted quantiles of the score distribution is
@@ V-B2 Conformal Prediction for Graphs
$\hat{C}(X)=\left\{Y\in\mathcal{Y}:s(X,Y)\leq Q_{1-\alpha}\left(\sum_{i=1}^{n}w_{i}\delta_{s_{i}}\right)\right\},$ (31)
where $Q_{1-\alpha}$, as defined in Eq. (28), is the $(1-\alpha)$-quantile of a distribution. The inductive cases are studied in [125], where the principle of exchangeability is not applicable. Based on the adaptive prediction set (APS) [112], they integrate non-exchangeable conformal prediction with the concept of homophily in graphs, leading to three distinct variants of the neighborhood adaptive prediction sets (NAPS) for the prediction set construction. The primary variant, NAPS, assigns the weights in Eq. (31) to $w_{i}=1$ if $X_{i}\in\mathcal{N}_{n+1}^{K}$, where $\mathcal{N}_{n+1}^{K}$ includes all nodes within $K$-hop neighborhood of node $X_{n+1}$. The second one, NAPS-H, adopts a hyperbolic decay rate for $w_{i}=k^{-1}$ where $k$ is the $k$-hop neighbor of node $X_{n+1}$. The third one, NAPS-G, uses a geometric decay rate for $w_{i}=2^{-k}$.
V-C Calibration
V-C1 Calibration Overview
[…]
Calibration Formulation. We begin with a classifier $f$ trained by a feature set and a corresponding label set whose label space is $\mathcal{Y}=\{1,...,K\}$. With inputs $x_{i}$ and the true label $y_{i}$ as an one-hot vector $\textbf{y}_{i}=(y_{i}^{1},...,y_{i}^{K})$ for $i=1,...,n$, $f$ applies softmax operation to output a probability vector, $\textbf{v}_{i}=(v_{i}^{1},...,v_{i}^{K}),\sum_{j=1}^{K}v_{i}^{j}=1$. The predicted label and confidence of $x_{i}$, denoted by $\hat{y}_{i}$ and $\hat{p}_{i}$ respectively, are
$\hat{y}_{i}=\arg\max_{j\in\mathcal{Y}}v_{i}^{j},\text{ }\hat{p}_{i}=\max_{j\in\mathcal{Y}}v_{i}^{j}.$ (32)
A model is well-calibrated if the probability of correct prediction equals the prediction confidence for all labels and all confidence levels [9], or more formally, $\text{Pr}(y_{i}=j|v_{i}^{j}=p)=p,\text{ }\forall j\in\mathcal{Y},\text{ }\forall p\in[0,1]$.
Calibration Method. Calibration can be conducted on top of accuracy-driven models, called post-processed methods. Histogram binning is initially proposed in [126] as a non-parametric calibration method. For a binary classification task with label space $\mathcal{Y}=\{0,1\}$ and $\hat{p}_{i,1}$ as the confidence of input $x_{i}$ in label 1, $B_{m}$ is a set including samples whose $\hat{p}_{i,1}$ is in the interval $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$ based on user-defined $M$, so we can use a calibration set to minimize
where $\theta_{m}$ is the estimated number of positive-class samples in the $B_{m}$ set. For a test sample $x_{n+1}$ whose $\hat{p}_{n+1,1}$ belong to $I_{m}$, $\hat{p}_{n+1,1}$ should be replaced by $\theta_{m}$ as a calibrated confidence. As Isotonic regression in [127] suggested, the number and size of $I_{m}$ can also be optimized. [128] applied isotonic regression to deep learning models for regression tasks with promising performance. Both histogram binning and isotonic regression can be applied to multi-class classification tasks using the one-vs-rest strategy. Platt scaling is a parametric calibration method introduced in [129]. Traditionally, $\textbf{v}_{i}$ is transformed from a logit vector $\textbf{z}_{i}$ via softmax function $\sigma(\cdot)$. Platt scaling generalizes the process by $\textbf{v}_{i}=\sigma(\textbf{W}\cdot\textbf{z}_{i}+\textbf{b})$, which can be optimized by minimizing negative log-likelihood on a calibration set. When W is a scaler and $\textbf{b}=0$, it is also referred to as temperature scaling.
Calibration can also be conducted during the training phase, without requiring an additional calibration dataset. Traditionally, we train a model with cross-entropy (CE), which is equivalent to NLL loss when we use a softmax output layer in a multi-class classification problem and encode labels as one-hot vectors. However, CE and NLL are minimized when a model assigns a high probability to the correct class for each instance. In other words, a model is rewarded for being overconfident in its correct predictions, and miscalibration can be caused by the overfitting of CE loss and NLL loss [130, 131]. To mitigate the issue, label smoothing tunes a parameter $\alpha\in[0,1]$, so the one-hot vector $\textbf{y}_{i}=(y_{i}^{1},...,y_{i}^{K})$ becomes smoother by
$\textbf{y}_{i}\rightarrow\textbf{y'}_{i}=\left((1-\alpha)y_{i}^{1}+\frac{\alpha}{K},...,(1-\alpha)y_{i}^{K}+\frac{\alpha}{K}\right),$ (34)
so it penalizes the model for assigning a full probability to the ground truth label [132]. Similarly, focal loss (FL) is introduced in [133] and it is applied to calibration in [131] with the loss function:
$\text{FL}(x_{i})=-(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i}),\gamma\geq 0.$ (35)
Focal loss only considers confidence in the true label and the loss value will decrease faster as we increase $\gamma$, so it can prevent being over-confident. If $\gamma=0$, FL goes back to CE.
V-C2 Calibration for Graphs
Accuracy alone is not enough for high-stake decision-making. The calibration performance of modern neural networks has been proven poor along a series of factors, including network width, depth, batch normalization, and weight decay [9]. Calibration on GNNs also started to draw people’s attention recently as well. GNNs for classification tasks can generate softmax outputs that can be regarded as probabilities that the predicted results are correct.
Node-level Classification. Traditional calibration techniques are typically applied to models handling independent and identically distributed (i.i.d.) data. In contrast, graph data, or relational data, presents a unique challenge due to its dependent nature. In graph data, the characteristics or status of one node may be influenced by the attributes or conditions of neighboring nodes, highlighting an inherent dependency where a node’s properties are not solely its own but also a reflection of its position within the graph [134]. This interdependency complicates the calibration process for GNNs. Furthermore, issues with class imbalance can lead to collapsed predictions, necessitating the adjustment of class weights during the training phase.
The calibration performance of GNNs on node-level classification tasks is related to GNNs’ architectures, neighbors’ labels, and the number of training epochs. The authors of [135, 136] validate the general tendency towards underconfidence for predictions of GNNs. Besides, it is proved that GNNs with wider layers tend to be better calibrated, and GNNs change from being under-confident to overconfident as the number of network layers increases  [137, 138]. Calibration error on the classification of a node can be influenced by the node’s neighbor as well. Experimental results of [135, 136] show that calibration error decreases and a node tends to have similar confidences with its neighbors when more neighbors have the same label as the center one. However, the results of  [137] reveal that well-calibration is more likely to occur when the same-class-neighbor ratio is around 0.4 to 0.5. Even though both results are drawn from GCN with the same datasets, the conflicts are probably caused by the different network structures and numbers of training epochs. As the experiments in  [137] demonstrate, GNNs tend to be under-confident with early stopping and will be overconfident with too many training epochs, as they start to overfit training data.
Conventional calibration methods can not recalibrate GNNs well. Compared with Monte-Carlo dropout (MC-dropout) [49], histogram binning [126], and isotonic regression [127], temperature scaling (TS) [129, 139] usually outputs the best calibration results in the experiment of  [134]. Similarly, experiments in [140] validate that the overall performance of Platt scaling (PS), TS, Accuracy-versus-Uncertainty (AvU) [141], and Graph Calibration Loss (GCL) [138], which will be introduced soon, is better than MC-dropout and Model Ensemble [142] techniques. Since MC-dropout and Model Ensemble rely on averaging predictions from multiple models, the limited architectural diversity of GCNs may cause their underperformance. Moreover, the authors of [140] find out models optimized on multi-class node classification tasks will be better calibrated and more discriminative than those trained on binary classification tasks.
The calibration methods for GNNs can be post-processed. CaGCN is a topology-aware post-hoc calibration GCN model developed in [135] assuming that neighboring nodes share similar ground truth confidence levels with the center one. With an undirected attributed GCN optimized for prediction accuracy, CaGCN works as another GCN to calibrate the outputs of the accuracy-driven model, thus it can be regarded as a Platt scaling method adapted to graphic models. To preserve the accuracy of the original classification GCN, the optimization objective function of CaGCN is a combination of negative log-likelihood loss and a customized calibration loss on validation nodes with a balancing hyperparameter. Graph Attention Temperature Scaling (GATS) [136] is also a post hoc method, and embeds the attention mechanism for node-wise calibration. For each node, a scaling process will consider global calibration bias, relative confidence compared with neighbors, and aggregated neighborhood influence with corresponding attention coefficients. Ratio-binned scaling (RBS), proposed in [137], follows the logic of histogram binning to calibrate outputs of a GNN. RBS will first estimate the same-class-neighbor ratio of each node, group them into $M$ bins, and finally learn a temperature for each bin to re-calibrate them.
Conducting calibration during model training is also attractive since it does not require additional data and computation costs. Graph Calibration Loss (GCL) is introduced in [138] as an end-to-end calibration method for GNNs. GCL is designed to minimize the discrepancy between the ground-truth probability distribution $p$ and predicted confidence distribution $\hat{p}$. Specifically, GCL is the KL divergence between $p$ and $\hat{p}$ regularized by the product of their entropy. HyperU-GCN is proposed in  [143] to calibrate the hyperparameter uncertainty of automated GNNs by defining it as a bi-level problem. The upper-level problem explains uncertainties by developing probabilistic hypernetworks through a variational Bayesian method. The lower-level problem learns how hyperparameter distribution is propagated to GCN weights.
To make GNNs robust to adversarial attacks and discriminatory biases, FairGNN [144], RobustGNN [145], and NIFTY [146] are proposed to enhance the stability of GNNs. However, these methods will make GNNs either overconfident in incorrect predictions or biased to global graph noises. To address the problem, Multi-view Confidence-calibrated NIFTY (MCCNIFTY) is introduced in [147]. Based on evidential theory, a novel node embedding learning module is developed. It includes an intra-view evidence calibration, an inter-view evidence fusion, and an uncertainty-aware message-passing process, optimizing for counterfactual fairness and stability at the sub-graph level.
Link Prediction. An important downstream task for GNN node classification is link prediction [148]. The authors of [149] study the calibration performance of link prediction and find that GNNs tend to be overconfident in negative link predictions but underconfident in positive ones. To address the issue, IN-N-OUT is proposed to parameterize a temperature for calibration. Two embeddings of an edge will be obtained when training the GNNs with and without the edge, and calculating the discrepancy between the two embeddings of the edge will help determine the value of temperature scaling.
Anchoring technique, named as $\Delta$-UQ, is introduced in [150] to estimate the uncertainty of predictions. During training, an anchor $c$ is randomly chosen for stochastic centering, and a relative representation $[x-c,c]$ of each input $x$ is created. The anchor value is different in each training iteration. For inference, $\Delta$-UQ collects multiple predictions of an input $x$ with different predefined $c$ and estimates uncertainty based on aggregated predictions.
The anchoring technique is adopted in [151] to calibrate the prediction uncertainty for link prediction. It defines link prediction by two processes: passing node information with a multi-layer deterministic encoder and predicting the label of two nodes with a stochastic decoder. The proposed method is named as E-$\Delta$UQ with three variants. The first variant applies anchoring between encoding and decoding, thus it is based on deterministic node features. The second variant implements one more anchoring after specific layers of encoding. Based on the second one, the last variant pre-trains the encoder with an auxiliary attribute masking task on node features.
Graph-level Classification. The anchoring technique is also adopted in [152], named as G-$\Delta$UQ with three variants, for the calibration of GNNs on graph-level classification tasks under the conditions of covariate and concept shifts using Superpixel-MNIST datasets [153, 154, 155, 156]. The first variant is node feature anchoring, implemented when inputting training data. A Gaussian distribution is fitted to node features in training data and an anchor is randomly selected from the distribution for each node. The second variant is intermediate anchoring, which applies stochastic centering for each node after specified hidden layers. Readout anchoring is the last variant applying anchoring to the final graph representation before a graph-level readout function, instead of anchoring for each node.

### v3

@@ V-D Conformal Prediction
Inductive CP, or split conformal prediction (SCP), is a widely applied version of CP [106]. For a trained classifier ${f}$ and input $x$ with label space $\mathcal{Y}=\{1,...,K\}$, the estimated probability for each class is $\textbf{v}=(v^{1},...,v^{K})$. Let $f(x)_{y}=v^{y}$ denote the predicted probability of the true class $y$ from the softmax layer. The calibration set $D_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is applied to quantify how uncertain the model $f$ would be. The score function $s(x,y)=1-f(x)_{y}$ is widely selected for classification tasks, and it measures how predictions conform to the true values at inputs. The conformal score for $(X_{i},Y_{i})\in D_{cal}$ is defined as $s_{i}=1-f(X_{i})_{Y_{i}}$. Let $S=\{s_{i}\}_{i=1}^{n}$, we define $\tau$ as the empirical $1-\alpha$ quantile of $S$. In other words, $\tau=\lceil(1-\alpha)(n+1)\rceil$-th smallest element of $S$. Formally, defining $\delta_{s_{i}}$ as a point mass at $s_{i}$, the quantile is
$\tau=\text{Quantile}\left(1-\alpha,\frac{1}{n}\sum\nolimits_{i=1}^{n}\delta_{s_{i}}\right).$ (39)
@@
For a test input $X_{n+1}$, a prediction set ${C}(X_{n+1})=\{k|1-f(X_{n+1})_{k}\leq\tau,k\in\mathcal{Y}\}$ includes all classes $k\in\mathcal{Y}=\{1,...,K\}$ whose conformal score is less or equal to $\tau$. If the test sample $(X_{n+1},Y_{n+1})$ is exchangeable with the elements of $D_{cal}$, a coverage guarantee holds that
@@
Transductive CP, or full conformal prediction (FCP), is one of the earliest approaches to CP [105]. With training data $D_{train}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a test input $X_{n+1}$, the trained model $f^{k}$ is based on $D_{train}\cup\{(X_{n+1},k)\},\forall k\in\{1,2,...,K\}$, so multiple models need to be trained. For each $f^{k}$, the conformal scores are $s_{i}^{k}=1-f^{k}(X_{i})_{Y_{i}}$ and $s_{n+1}^{k}=1-f^{k}(X_{n+1})_{k}$. If $s_{n+1}^{k}$ is less than the $1-\alpha$ quantile of $\{s_{i}^{k}\}_{i=1}^{n+1}$, we put the corresponding $k$ to the prediction set of $X_{n+1}$.
@@
To make the size of $C(X_{n+1})$ more adaptive to $X_{n+1}$ under the exchangeability assumption, novel conformal score functions are proposed  [109, 209, 108, 210]. However, the exchangeability assumption can be violated by distribution shift between calibration distribution $P_{XY}$ and test distribution $Q_{XY}$ such that $(X_{n+1},Y_{n+1})$ is non-exchangeable with the elements of $D_{cal}$. In the case of covariate shift ($P_{X}\neq Q_{X}$), an importance weighting technique is introduced based on likelihood ratio to maintain the coverage guarantee in Eq. (40[211]. When joint distribution shift ($P_{X}\neq Q_{X}$, $P_{Y|X}\neq Q_{Y|X}$) occurs, prediction sets can be modified to keep $1-\alpha$ coverage on test data under dynamic distribution shift [212, 213], and static distribution shift [110, 214, 215]. It is proved that the coverage gap can be bounded by the total variation distance [216] and Wasserstein distance [217] between calibration and test conformal scores under the non-exchangeable condition.
Conformal Prediction for Graphs. When labeled calibration nodes and test nodes are randomly distributed throughout the graph, it is reasonable to assume that calibration and test nodes are exchangeable. Under this setting, marginal coverage guarantees remain valid. JuryGCN [111] estimates prediction uncertainty in node classification tasks using jackknife methods. In graph settings, node homophily plays an important role: if a node is surrounded by neighbors with similar labels, their features provide strong predictive cues, resulting in higher model confidence. Conversely, when a node is surrounded by neighbors with different labels (i.e., heterophily), the uncertainty in prediction should increase. Motivated by this observation, Diffused Adaptive Prediction Sets (DAPS) [112] leverage the graph structure via neighborhood diffusion and define the diffused score as:
$s^{\text{diff}}(X_{i},Y_{i})=(1-\beta)s(X_{i},Y_{i})+\frac{\beta}{|\mathcal{N}_{i}|}\sum{X_{j}\in\mathcal{N}_{i}}s(X_{j},Y_{i}),$
where $\beta$ signifies the diffusion parameter and $\mathcal{N}_{i}$ includes the features of node $i$ neighbors. The diffused score function $s^{\text{diff}}$ preserves exchangeability if the original scores $s(X_{i},Y_{i})$ are exchangeable. This diffusion approach is particularly beneficial in homophilic graphs, where neighboring nodes tend to share similar ground-truth labels. In such settings, the prediction for a node can be improved by incorporating the information from its structurally close, and often label-consistent, neighbors.
However, when calibration and test data correspond to different subgraphs, such as disjoint node sets with limited connectivity, the exchangeability assumption typically breaks down due to structural and distributional differences between subgraphs. This is because the calibration nodes and test nodes are no longer drawn from the same distributional context within the graph, especially if they differ in topological properties, feature distributions, or label characteristics. For example, if the calibration nodes all lie within a densely connected community or are concentrated in a homophilic region of the graph, while the test nodes are located in sparser or more heterophilic regions, the conformity scores computed on the calibration set may not generalize to the test nodes. This mismatch undermines the core assumption of exchangeability required for valid coverage guarantees. Furthermore, the violation of exchangeability is exacerbated in inductive settings where the test node influences (or is influenced by) calibration nodes through shared neighbors or overlapping receptive fields in message-passing architectures. This issue has been highlighted in recent works such as [112], which emphasize that topological locality and data partitioning within graphs introduce structured dependencies that vanilla CP methods are not designed to handle. Beyond the standard assumption of exchangeability, Barber et al. [216] propose a generalization of conformal prediction known as Non-exchangeable Conformal Prediction (NexCP), which accommodates settings where data points are not identically distributed. Instead of treating all calibration samples equally, NexCP assigns deterministic weights $w_{1},\dots,w_{n}\in[0,1]$ to $n$ calibration samples, and generates the prediction set as
${C}(X_{n+1})=\{k$ $|s(X_{n+1},k)\leq$ (41)
@@
$\text{Quantile}\left(1-\alpha,\sum\nolimits_{i=1}^{n}w_{i}\delta_{s_{i}}\right),k\in\mathcal{Y}\}.$
The core challenge lies in selecting appropriate weights $w_{i}$ that reflect the similarity between calibration samples and the test input $X_{n+1}$. Intuitively, calibration examples that come from distributions closer to that of the test sample should receive higher weights, as they are more informative and relevant for estimating uncertainty. This idea has been extended to graph data by Clarkson et al.[113], who integrate NexCP into the Adaptive Prediction Sets (APS) framework [109]. In graphs, homophily is a common structural property. Leveraging this, Neighborhood Adaptive Prediction Sets (NAPS) are proposed to assign weights to calibration nodes based on their distances to the test node. Nodes that are closer (in terms of hop count or diffusion distance) are considered more exchangeable and are thus given greater influence in determining the prediction set. Three variants of NAPS have been developed, each capturing different notions of neighborhood similarity and diffusion dynamics within the graph. The primary variant, NAPS, assigns the weights $w_{i}=1$ if $X_{i}\in\mathcal{N}_{n+1}^{K}$, where $\mathcal{N}_{n+1}^{K}$ includes all nodes within $K$-hop neighborhood of node $X_{n+1}$. The second one, NAPS-H, adopts a hyperbolic decay rate for $w_{i}=k^{-1}$ where $k$ is the $k$-hop neighbor of node $X_{n+1}$. The third one, NAPS-G, uses a geometric decay rate for $w_{i}=2^{-k}$.
In addition to node classification, conformal prediction has been applied to other graph-related tasks, such as link prediction [218, 219, 220] and graph regression [221, 222]. However, these studies largely extend standard inductive CP to new task settings while still relying on the assumption of exchangeability. As a result, we do not elaborate on these works further.